# Optimal Transport for Measures with Noisy Tree Metric

### v1

@@ 4 Related work and discussion
Notice that the robust OT approach for OT problem with noisy ground cost is dissimilar to the Wasserstein distributionally robust optimization (Kuhn et al., 2019; Blanchet et al., 2021). Although they may share the min-max formulation, the Wasserstein distributionally robust optimization seeks the best data-driven decision under the most adverse distribution from a Wasserstein ball of a certain radius. Additionally, one should distinguish this robust OT approach for noisy ground cost with the outlier-robust approach for OT where the noise is on input probability measures (Balaji et al., 2020; Mukherjee et al., 2021; Le et al., 2021; Nietert et al., 2022).
Leveraging tree structure to scale up OT problems has been explored for standard OT (Le et al., 2019; Yamada et al., 2022), for OT problem with input measures having different total mass (Sato et al., 2020; Le & Nguyen, 2021), and for Wasserstein barycenter (Takezawa et al., 2022). To our knowledge, our work is the first approach to exploit tree structure over supports to scale up robust OT approach for OT problem with noisy ground cost. Furthermore, notice that max-sliced Wasserstein (Deshpande et al., 2019) is $1$-dimensional OT-based approach for the max-min robust OT. However, there are no fast/efficient algorithmic approaches yet due to its non-convexity. Our approach is based on tree structure which provides more flexibility and degrees of freedoms to capture the topological structure of input probability measures than the 1-dimensional OT-based approach (i.e., choosing a tree rather than a line). Moreover, our novel uncertainty sets of tree metrics play the key role to scale up the computation of robust OT. The uncertainty sets not only includes a diversity of tree structures, but also a variety of edge lengths in an elegant framework following the theoretical guidance in Theorem 3.1.
@@ 5 Experiments
Document classification. We evaluate on $4$ real-world document datasets: TWITTER, RECIPE, CLASSIC, and AMAZON [^7] Although these document datasets may be noisy (Sato et al., 2022), we have no assumption about the cleanliness for datasets used in our experiments. Therefore, our experiments on these datasets can be regarded as evaluating OT problem for measures with noisy tree metric in the same noisy dataset settings.. Their characteristics are listed in Figure 2. We follow the same approach in (Le et al., 2019) to embed words into vectors in $\mathbb{R}^{300}$, and represent each document as a probability measure where its supports are in $\mathbb{R}^{300}$.
@@
TDA. We consider orbit recognition on Orbit dataset and object shape recognition on MPEG7 dataset for TDA. We summarize the characteristics of these datasets in Figure 3. We follow the same approach in (Le et al., 2019) to extract persistence diagram (PD) for orbits and object shapes, which are multisets of points in $\mathbb{R}^{2}$, and represent each PD as a probability measure where its supports are in $\mathbb{R}^{2}$.
Noisy tree metric. We apply the clustering-based tree metric sampling method (Le et al., 2019) to obtain a tree metric over supports. We then generate perturbations by deviating each tree edge length by a random nonnegative amount which is less than or equal $\Delta\in\mathbb{R}_{+}$ (i.e., $|w_{e}-w_{e}^{*}|\leq\Delta$) where $w_{e}^{*},w_{e}$ are tree edge lengths on the tree before and after the perturbations respectively. We set $\Delta=0.5$ for document classification (for edge lengths constructed from supports in $\mathbb{R}^{300}$); and set $\Delta=0.05$ for TDA (for edge lengths constructed from supports in $\mathbb{R}^{2}$).
@@
Experimental setup. We apply kernel SVM for the proposed robust OT kernels $k_{\textsf{RT}_{{\mathcal{U}}}}$ and $k_{\textsf{RT}_{{\mathcal{U}}_{p}}}$ and the counterpart standard OT (i.e., TW) kernel $k_{\textsf{TW}}$ for measures with a given noisy tree metric on document classification and TDA. For $\textsf{RT}_{{\mathcal{U}}}$, we consider $\alpha=\min(\lambda\mathbbm{1},w)$ and $\beta=\lambda\mathbbm{1}$ where minimum operator is element-wise; $\lambda$ is the radius in $\textsf{RT}_{{\mathcal{U}}_{p}}$; and recall that $w$ is a vector of edge lengths of the given tree. We set $p=2$ for the robust OT $\textsf{RT}_{{\mathcal{U}}_{p}}$ (or $\textsf{RT}_{{\mathcal{U}}_{2}}$).
@@
Empirical results and discussion. Figures 2 and 3 illustrate the performances of SVM for document classification and TDA respectively. The performances of the proposed kernels $k_{\textsf{RT}_{{\mathcal{U}}}}$ and $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$ compare favorably to those of the counterpart OT kernel $k_{\textsf{TW}}$. Notably, kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$ improves about 15% average accuracy over kernel $k_{\textsf{TW}}$ on Orbit. In addition, kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$ consistently outperforms kernel $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$, except on MPEG7 where their performances are comparable, which may come from the freedom to constraint over each edge length in ${\mathcal{U}}$ of kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$. Moreover, our results also agree with previous observations on other existing max-min/min-max robust OT where the robust OT approach improves performances of the counterpart standard OT for measures with noisy ground cost. For further discussions, we refer the readers to the supplementary (§C–§D).
@@
Effects of the radius $\lambda$. Figures 4 and 5 illustrate the effects of the radius $\lambda$ on the proposed robust OT kernels on document classification and TDA respectively. Notice that when $\lambda=0$, the max-min robust OT for are equivalent to the counterpart standard OT. We observe that kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$ is less sensitive with the radius $\lambda$ than kernel $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$. The performances of kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$ gradually increase when the radius $\lambda$ increases, after these performances reach their peaks, they decrease when $\lambda$ increases. The performances of kernel $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$ also share a similar pattern but more noisy. Therefore, cross validation for the radius $\lambda$ is useful in applications, especially for kernel $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$ in our simulations.
@@ 6 Conclusion
In this work, we proposed novel uncertainty sets of tree metrics which not only include metric metrics with varying edge lengths, but also having diverse tree structures in an elegant framework. By building upon these uncertainty sets and leveraging tree structure, we scale up the max-min robust OT approach for OT problem for probability measures with noisy tree metric. Moreover, by exploiting the negative definiteness of the robust OT, we proposed positive definite kernels built upon the robust OT and evaluated them for kernel SVM on document classification and TDA. For future work, extending the problem settings for more general applications (e.g., by leveraging the clustering-based tree metric sampling method), or for more general structures (e.g., graphs) is an interesting direction.
@@ A.1 Kernels
Positive Definite Kernels (Berg et al., 1984, pp. 66–67). A kernel function $k:\Omega\times\Omega\rightarrow\mathbb{R}$ is positive definite if for every positive integer $m\geq 2$ and every points $x_{1},x_{2},...,x_{m}\in\Omega$, we have
@@
Negative Definite Kernels (Berg et al., 1984, pp. 66–67). A kernel function $k:\Omega\times\Omega\rightarrow\mathbb{R}$ is negative definite if for every integer $m\geq 2$ and every points $x_{1},x_{2},...,x_{m}\in\Omega$, we have
@@
Theorem 3.2.2 in (Berg et al., 1984, pp. 74). Let $\kappa$ be a negative definite kernel function. Then, for every $t>0$, the kernel
@@
Definition 2.6 in (Berg et al., 1984, pp. 76). A positive definite kernel $\kappa$ is infinitely divisible if for each $n\in{\mathbb{N}}^{*}$, there exists a positive definite kernel $\kappa_{n}$ such that
@@
Corollary 2.10 in (Berg et al., 1984, pp. 78). Let $\kappa$ be a negative definite kernel function. Then, for $10<t<$ 1, the kernel
@@ A.2 Tree Metric
Tree metric. A metric $d:\Omega\times\Omega\rightarrow\mathbf{R}$ is a tree metric on $\Omega$ if there exists a tree ${\mathcal{T}}$ with non-negative edge lengths such that all elements of $\Omega$ are contained in its nodes and such that for every $x,z\in\Omega$, we have $d(x,z)$ equals to the length of the path between $x$ and $z$ (Semple & Steel, 2003, §7, pp.145–182). We write $d_{{\mathcal{T}}}$ for the tree metric corresponding to the tree ${\mathcal{T}}$.
Clustering-based tree metric sampling method. The clustering-based tree metric sampling method was proposed by Le et al. (2019) (see their §4). Le et al. (2019) also reviewed the farthest-point clustering in §4.2 in their supplementary, which is the main component used in the clustering-based tree metric sampling method.
@@ A.3 Optimal Transport (OT) for Measures on a Tree
We refer the readers to (Kusano et al., 2017, §2) for a brief review about the mathematical framework for persistence diagrams (e.g., persistence diagrams, filtrations, persistent homology).
@@ Appendix B Proofs
and suppose that $\alpha=\beta=\lambda\mathbbm{1}$, it follows from the definition of the $\ell_{\infty}$-norm that ${\mathcal{U}}({\mathcal{T}},\alpha,\beta)=\widetilde{\mathcal{U}}_{\infty}({\mathcal{T}},\lambda)$. Thus,
@@
Thanks to formula (7) for $\textsf{RT}_{{\mathcal{U}}({\mathcal{T}},\alpha,\beta)}$ which is independent of $\alpha$, we can further drop the condition $\alpha=\lambda\mathbbm{1}$. That is, connection (21) holds true under the only condition $\beta=\lambda\mathbbm{1}$. Thus, the proof is completed.
@@ C.1 Related Work
Yamada et al. (2022) considered to use TW as an approximation model for some given (oracle) $1$-Wasserstein distance. In particular, given some supervised triplets $(\mu,\nu,{\mathcal{W}}(\mu,\nu))$ with measures $\mu,\nu$ being supported on high-dimensional vector space $\mathbb{R}^{n}$. The ground cost metric for the observed $1$-Wasserstein is unknown. Yamada et al. (2022) used TW as a model to fit the observed $1$-Wasserstein distances, i.e., estimate tree metric such that ${\mathcal{W}}_{{\mathcal{T}}}(\cdot,\cdot)={\mathcal{W}}(\cdot,\cdot)$. Therefore, the approach in (Yamada et al., 2022) cannot be used in our considered problem due to the following two main reasons:
@@
  • (ii) our considered probability measures are supported in a given tree metric space $({\mathcal{T}},d_{{\mathcal{T}}})$. Moreover, the given tree structure is not necessarily a physical tree in the sense that the set of vertices $V$ is a subset of some high-dimensional vector space $\mathbb{R}^{n}$, and each edge $e$ is the standard line segment in $\mathbb{R}^{n}$ connecting the two end-points of edge $e$. The given tree structure ${\mathcal{T}}$ in our problem can be non-physical, while in (Yamada et al., 2022), input probability measures $\mu,\nu$ are supported on some high-dimensional vector space $\mathbb{R}^{n}$.

### v2

@@ 4 RELATED WORK AND DISCUSSION
Notice that the robust OT approach for OT problem with noisy ground cost is dissimilar to the Wasserstein distributionally robust optimization (Kuhn et al., 2019; Blanchet et al., 2021). Although they may share the min-max formulation, the Wasserstein distributionally robust optimization seeks the best data-driven decision under the most adverse distribution from a Wasserstein ball of a certain radius. Additionally, one should distinguish this robust OT approach for noisy ground cost with the outlier-robust approach for OT where the noise is on input probability measures (Balaji et al., 2020; Mukherjee et al., 2021; Nguyen et al., 2021a; Nietert et al., 2022).
Leveraging tree structure to scale up OT problems has been explored for standard OT (Le et al., 2019; Yamada et al., 2022), for OT problem with input measures having different total mass (Sato et al., 2020; Le & Nguyen, 2021), and for Wasserstein barycenter (Takezawa et al., 2022). To our knowledge, our work is the first approach to exploit tree structure over supports to scale up robust OT approach for OT problem with noisy ground cost. Furthermore, notice that max-sliced Wasserstein (Deshpande et al., 2019) is one-dimensional OT-based approach for the max-min robust OT. However, there are no fast/efficient algorithmic approaches yet due to its non-convexity. Our approach is based on tree structure which provides more flexibility and degrees of freedoms to capture the topological structure of input probability measures than the one-dimensional OT-based approach (i.e., choosing a tree rather than a line). Moreover, our novel uncertainty sets of tree metrics play the key role to scale up the computation of robust OT. The uncertainty sets not only includes a diversity of tree structures, but also a variety of edge lengths in an elegant framework following the theoretical guidance in Theorem 3.1.
@@ 5 EXPERIMENTS
Document classification.
We evaluate on $4$ real-world document datasets: TWITTER, RECIPE, CLASSIC, and AMAZON. Their characteristics are listed in Figure 2. We follow the same approach in (Le et al., 2019) to embed words into vectors in $\mathbb{R}^{300}$, and represent each document as a probability measure where its supports are in $\mathbb{R}^{300}$.
@@
TDA.
We consider orbit recognition on Orbit dataset and object shape recognition on MPEG7 dataset for TDA. We summarize the characteristics of these datasets in Figure 3. We follow the same approach in (Le et al., 2019) to extract persistence diagram (PD) for orbits and object shapes, which are multisets of points in $\mathbb{R}^{2}$, and represent each PD as a probability measure where its supports are in $\mathbb{R}^{2}$.
Noisy tree metric.
We apply the clustering-based tree metric sampling method (Le et al., 2019) to obtain a tree metric over supports. We then generate perturbations by deviating each tree edge length by a random nonnegative amount which is less than or equal $\Delta\in\mathbb{R}_{+}$ (i.e., $|w_{e}-w_{e}^{*}|\leq\Delta$) where $w_{e}^{*},w_{e}$ are tree edge lengths on the tree before and after the perturbations respectively. We set $\Delta=0.5$ for document classification (for edge lengths constructed from supports in $\mathbb{R}^{300}$); and set $\Delta=0.05$ for TDA (for edge lengths constructed from supports in $\mathbb{R}^{2}$).
@@
Experimental setup.
We apply kernel SVM for the proposed robust OT kernels $k_{\textsf{RT}_{{\mathcal{U}}}}$ and $k_{\textsf{RT}_{{\mathcal{U}}_{p}}}$ and the counterpart standard OT (i.e., TW) kernel $k_{\textsf{TW}}$ for measures with a given noisy tree metric on document classification and TDA. For $\textsf{RT}_{{\mathcal{U}}}$, we consider $\alpha=\min(\lambda\mathds{1},w)$ and $\beta=\lambda\mathds{1}$ where minimum operator is element-wise; $\lambda$ is the radius in $\textsf{RT}_{{\mathcal{U}}_{p}}$; and recall that $w$ is a vector of edge lengths of the given tree. We set $p=2$ for the robust OT $\textsf{RT}_{{\mathcal{U}}_{p}}$ (or $\textsf{RT}_{{\mathcal{U}}_{2}}$).
@@
Empirical results and discussion.
Figures 2 and 3 illustrate the performances of SVM for document classification and TDA respectively. The performances of the proposed kernels $k_{\textsf{RT}_{{\mathcal{U}}}}$ and $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$ compare favorably to those of the counterpart OT kernel $k_{\textsf{TW}}$. Notably, kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$ improves about 15% average accuracy over kernel $k_{\textsf{TW}}$ on Orbit. In addition, kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$ consistently outperforms kernel $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$, except on MPEG7 where their performances are comparable, which may come from the freedom to constraint over each edge length in ${\mathcal{U}}$ of kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$. Moreover, our results also agree with previous observations on other existing max-min/min-max robust OT where the robust OT approach improves performances of the counterpart standard OT for measures with noisy ground cost. For further discussions, we refer the readers to the supplementary (§C–§D).
@@
Effects of the radius $\lambda$.
Figures 4 and 5 illustrate the effects of the radius $\lambda$ on the proposed robust OT kernels on document classification and TDA respectively. Notice that when $\lambda=0$, the max-min robust OT for are equivalent to the counterpart standard OT. We observe that kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$ is less sensitive with the radius $\lambda$ than kernel $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$. The performances of kernel $k_{\textsf{RT}_{{\mathcal{U}}}}$ gradually increase when the radius $\lambda$ increases, after these performances reach their peaks, they decrease when $\lambda$ increases. The performances of kernel $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$ also share a similar pattern but more noisy. Therefore, cross validation for the radius $\lambda$ is useful in applications, especially for kernel $k_{\textsf{RT}_{{\mathcal{U}}_{2}}}$ in our simulations.
@@ 6 CONCLUSION
In this work, we proposed novel uncertainty sets of tree metrics which not only include metric metrics with varying edge lengths, but also having diverse tree structures in an elegant framework. By building upon these uncertainty sets and leveraging tree structure, we scale up the max-min robust OT approach for OT problem for probability measures with noisy tree metric. Moreover, by exploiting the negative definiteness of the robust OT, we proposed positive definite kernels built upon the robust OT and evaluated them for kernel SVM on document classification and TDA. For future work, extending the problem settings for more general applications (e.g., by leveraging the clustering-based tree metric sampling method (Le et al., 2019)), and/or for more general structures (e.g., graphs as in the Sobolev transport (Le et al., 2022, 2023)) are interesting research directions.
Acknowledgements
We thank the area chairs and anonymous reviewers for their comments. KF has been supported in part by Grant-in-Aid for Transformative Research Areas (A) 22H05106. TL gratefully acknowledges the support of JSPS KAKENHI Grant number 23K11243.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Yes]
    4. (d)
      Information about consent from data providers/curators. [Yes]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Supplement to “Optimal Transport for Measures with Noisy Tree Metric”
In this supplementary, we give brief reviews about some aspects used in our work, e.g., kernels, tree metric, and optimal transport on for probability measures on a tree in §A. We present detailed proofs for the theoretical results in §B, and give additional discussions about our work in §C. Further experimental results are placed in §D.
@@ A.1 Kernels
Positive Definite Kernels (Berg et al., 1984, pp. 66–67).
A kernel function $k:\Omega\times\Omega\rightarrow\mathbb{R}$ is positive definite if for every positive integer $m\geq 2$ and every points $x_{1},x_{2},...,x_{m}\in\Omega$, we have
@@
Negative Definite Kernels (Berg et al., 1984, pp. 66–67).
A kernel function $k:\Omega\times\Omega\rightarrow\mathbb{R}$ is negative definite if for every integer $m\geq 2$ and every points $x_{1},x_{2},...,x_{m}\in\Omega$, we have
@@
Theorem 3.2.2 in Berg et al. (1984, pp. 74).
Let $\kappa$ be a negative definite kernel function. Then, for every $t>0$, the kernel
@@
Definition 2.6 in Berg et al. (1984, pp. 76).
A positive definite kernel $\kappa$ is infinitely divisible if for each $n\in{\mathbb{N}}^{*}$, there exists a positive definite kernel $\kappa_{n}$ such that
@@
Corollary 2.10 in Berg et al. (1984, pp. 78).
Let $\kappa$ be a negative definite kernel function. Then, for $10<t<$ 1, the kernel
@@ A.2 Tree Metric
Tree metric.
A metric $d:\Omega\times\Omega\rightarrow\mathbf{R}$ is a tree metric on $\Omega$ if there exists a tree ${\mathcal{T}}$ with non-negative edge lengths such that all elements of $\Omega$ are contained in its nodes and such that for every $x,z\in\Omega$, we have $d(x,z)$ equals to the length of the path between $x$ and $z$ (Semple & Steel, 2003, §7, pp.145–182). We write $d_{{\mathcal{T}}}$ for the tree metric corresponding to the tree ${\mathcal{T}}$.
Clustering-based tree metric sampling method.
The clustering-based tree metric sampling method was proposed by Le et al. (2019) (see their §4). Additionally, Le et al. (2019) reviewed the farthest-point clustering in §4.2 in their supplementary, which is the main component used in the clustering-based tree metric sampling method.
@@ A.3 Optimal Transport (OT) for Measures on a Tree
We refer the readers to Kusano et al. (2017, §2) for a brief review about the mathematical framework for persistence diagrams (e.g., persistence diagrams, filtrations, persistent homology).
@@ Appendix B PROOFS
and suppose that $\alpha=\beta=\lambda\mathds{1}$, it follows from the definition of the $\ell_{\infty}$-norm that ${\mathcal{U}}({\mathcal{T}},\alpha,\beta)=\widetilde{\mathcal{U}}_{\infty}({\mathcal{T}},\lambda)$. Thus,
@@
Thanks to formula (7) for $\textsf{RT}_{{\mathcal{U}}({\mathcal{T}},\alpha,\beta)}$ which is independent of $\alpha$, we can further drop the condition $\alpha=\lambda\mathds{1}$. That is, connection (21) holds true under the only condition $\beta=\lambda\mathds{1}$. Thus, the proof is completed.
@@ C.1 Related Work
Yamada et al. (2022) considered to use TW as an approximation model for some given (oracle) $1$-Wasserstein distance. In particular, given some supervised triplets $(\mu,\nu,{\mathcal{W}}(\mu,\nu))$ with measures $\mu,\nu$ being supported on high-dimensional vector space $\mathbb{R}^{n}$. The ground cost metric for the observed $1$-Wasserstein is unknown. Yamada et al. (2022) used TW as a model to fit the observed $1$-Wasserstein distances, i.e., estimate tree metric such that ${\mathcal{W}}_{{\mathcal{T}}}(\cdot,\cdot)={\mathcal{W}}(\cdot,\cdot)$. Therefore, the approach in Yamada et al. (2022) cannot be used in our considered problem due to the following two main reasons:
@@
  • (ii) our considered probability measures are supported in a given tree metric space $({\mathcal{T}},d_{{\mathcal{T}}})$. Moreover, the given tree structure is not necessarily a physical tree in the sense that the set of vertices $V$ is a subset of some high-dimensional vector space $\mathbb{R}^{n}$, and each edge $e$ is the standard line segment in $\mathbb{R}^{n}$ connecting the two end-points of edge $e$. The given tree structure ${\mathcal{T}}$ in our problem can be non-physical, while in Yamada et al. (2022), input probability measures $\mu,\nu$ are supported on some high-dimensional vector space $\mathbb{R}^{n}$.